OpenStudy (anonymous):

A is the midpoint of XY Find: XA Find: AY and XY X--^3x---A---^5x-6--Y

OpenStudy (luigi0210):

\[\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile } \]

OpenStudy (anonymous):

AHHAHAHA^

OpenStudy (jdoe0001):

w000t the smokes w45 dat?

OpenStudy (jdoe0001):

the dreaded clipboard =P

OpenStudy (anonymous):

\Huge{\color{purple}{\textbf{W}} \color{orange}{\cal{E}} \color{green}{\mathbb{L}} \color{blue}{\mathsf{C}} \color{maroon}{\rm{O}} \color{red}{\tt{M}} \color{gold}{\tt{E}} \space \color{orchid}{\mathbf{T}} \color{Navy}{\mathsf{O}} \space \color{OrangeRed}{\boldsymbol{O}} \color{Olive}{\mathbf{P}} \color{Lime}{\textbf{E}} \color{DarkOrchid}{\mathsf{N}} \color{Tan}{\mathtt{S}} \color{magenta}{\mathbb{T}} \color{goldenrod}{\mathsf{U}} \color{ForestGreen}{\textbf{D}} \color{Salmon}{\mathsf{Y}} \ddot \smile }

OpenStudy (jdoe0001):

hmm, you forgot to use $$ delimeters

OpenStudy (jdoe0001):

 notice that if A is the middle point, then segments on either side of A are of equal length, thus we could say that XA = AY or \(\bf 3x = 5x-6\) solve for "x"